A merchents has coffee worth $60 a pound that she wants to mix with 70 pounds of coffe worth $90 a pound to get a mixture that is worth $70 a pound. How many pounds (lbs) of the $60 coffee should be used?
Answer by Maths68(1474)

Coffee-A
========
Worth = $60/pound
Weight=x pounds

Coffee-B
========
Worth = $90/pound
Weight= 70 pounds


Mixture
========
Worth = $70/pound
Weight= (x+70) pounds

(Coffee-A worth)*(Coffee-A weight) + (Coffee-B worth)*(Coffee-B weight) = (Mixture worth)*(Mixture weight)
(60)(x)+(90)(70)=(70)(x+70)
60x+6300=70x+4900
60x-70x=4900-6300
-10x=-1400
Divide by -10 both sides of above equation
-10x/-10=-1400/10
x=140

She should mix 140 pounds of coffee worth $60 a pound with 70 pounds of coffee worth $90 a pound to get a mixture that is worth $70 a pound